# DEPTH: Discourse Education through Pre-Training Hierarchically

### v1

@@ Abstract
Language Models (LMs) often struggle with linguistic understanding at the discourse level, even though discourse patterns such as coherence, cohesion, and narrative flow are prevalent in their pre-training data. Current methods address these challenges only after the pre-training phase, relying on expensive human annotated data to align the model. To improve the discourse capabilities of LMs already at the pre-training stage, we introduce DEPTH, an encoder-decoder model that learns to represent sentences using a discourse-oriented pre-training objective. DEPTH combines hierarchical sentence representations with two objectives: (1) Sentence Un-Shuffling, and (2) Span-Corruption. This approach trains the model to represent both sub-word-level and sentence-level dependencies over a massive amount of unstructured text. When trained either from scratch or continuing from a pre-trained T5 checkpoint, DEPTH learns semantic and discourse-level representations faster than T5, outperforming it in span-corruption loss despite the additional sentence-un-shuffling objective. Evaluations on the GLUE, DiscoEval, and NI benchmarks demonstrate DEPTH’s ability to quickly learn diverse downstream tasks, which require syntactic, semantic, and discourse capabilities. Overall, our approach extends the discourse capabilities of T5, while minimally impacting other natural language understanding (NLU) capabilities in the resulting LM. [^1] We release our repository for pre-training and fine-tuning encoder-decoder models as well as the checkpoints we used for this study.
@@ 1 Introduction
Discourse understanding—the ability to understand how sentences and broader textual units form cohesive narratives—is fundamental to effective communication. However, LMs often struggle with this, especially when dealing with long and complex inputs. This limitation hinders the performance of LMs on various natural language processing (NLP) tasks such as persuasive argumentation [16], summarization [69], essay scoring [36], dialogue systems [20], and following intricate instructions [60]. Existing LMs frequently generate text that lacks compositionality, suffers from disconnected ideas and contradictions, or fails to develop core ideas across sentences [17, 34, 35]. These issues stem from their limited ability to capture the underlying structure, implied meanings, and relationships between ideas that are crucial for effective communication.
Early attempts to incorporate discourse awareness into pre-training, such as Next Sentence Prediction (NSP) in BERT [12] and Sentence Order Prediction (SOP) in ALBERT [26], proved overly simplistic and hindered learning effective discourse representations [32, 26, 45]. Subsequent models like Sentence-level Language Model (SLM) [27] and CONPONO [21] improved discourse capabilities in LMs but lacked generative capabilities. On the other hand, the next-token prediction pre-training objective provided LMs like Generative Pre-trained Transformer (GPT) [43, 44, 4, 40] with powerful generative capabilities. However, without a dedicated and costly alignment phase [42], these LMs tend to falter at discourse-level evaluation benchmarks, even when scaled to over 100 billion parameters [35, 34, 5]. We therefore focus our studies on encoder-decoder models such as Text-to-Text Transfer Transformer (T5) [45], which can compete with (and sometimes outperform) GPT-3 [4] on long (e.g., paragraph-level) discourse-oriented tasks even at $\approx\frac{1}{400}$ the size [35, 34]. [^2] The T5 model used in this study, released by Wei et al. [60], was fine-tuned to follow instructions.
@@
To improve the discourse-capabilities of encoder-decoder models, we propose DEPTH (Discourse-Education through Pre-Training Hierarchically), a hierarchical language model that learns representations for both sub-word and sentence-level tokens. DEPTH builds on the architecture of T5 and incorporates the discourse-oriented objective of SLM. By employing a hierarchical attention mechanism across sub-word and sentence level tokens, DEPTH captures both fine-grained semantic dependencies and broader inter-sentence relationships. Furthermore, its simplified loss function unifies a discourse-oriented pre-training objective, previously limited to encoders, with the traditional teacher-forcing optimization used in decoders. When randomly initialized, our DEPTH model obtains meaningful representations for downstream tasks much faster than our baseline T5. Continuously pre-training T5 models with the DEPTH objective improves the discourse capabilities of the resulting models, without sacrificing performance in downstream NLU tasks.
@@ 2 Method
Pre-training DEPTH involves simultaneously performing span corruption [45], while also un-shuffling sentences as in Lee et al. [27]. In Section 2.1 we introduce a new tokenizer for DEPTH, which combines T5’s tokenizer with the sentence segmentation operation required for sentence un-shuffling. In Section 2.2, we detail how we combined the pre-training objectives of both models into the sequence-to-sequence framework of T5. Next, in Section 2.3, we discuss how we introduce hierarchical representations during pre-training, and how this hierarchy encourages the model to learn discourse representations. Finally, in Section 2.4, we demonstrate how to combine the losses of T5 and SLM into a unified objective that is conducive to traditional teacher-forcing.
@@ 2.1 Tokenization
We introduce a tokenization function, $t$, which transforms an input string, $s$, into a sequence of tokens, $X=(x_{1,1},x_{1,2},\ldots,x_{m,len(m)})$, in our vocabulary, $V$. In this formulation, each token $x_{i,j}$ denotes the $j$’th token of the $i$’th sentence, $m$ represents the number of sentences, and $len(i)$ represents the length of the $i$’th sentence. $V$ also includes special tokens <EOS>, <BOS>, <PAD>, and a set of sentinel tokens, $V_{sentinel}$, of the form <special_token_z> as in the original T5 paper.
Figure 1: Given an input document, like “Hello! How are you?”, DEPTH introduces sentence tokens (<SENT_i> and <EOSEN>) as part of tokenization. Then, as part of corruption, DEPTH applies both span masking (in this case, applying a single masked span over the tokens “are” and “you?”), and shuffles the sentences in the encoder with probability $0.5$. As part of training, sentence tokens in the encoder can only attend to tokens in their own sentence. Sentence tokens in the decoder can only attend to earlier sentence tokens (both in the encoder and in the decoder). Attention among sentence tokens is represented as arrows, where the source is the query token, and the target is the key token (cross attention uses dotted arrows, while self attention uses full arrows).
To facilitate DEPTH’s hierarchical structure, we segment sentences using NLTK [3] (as in SLM) in each example, and create $k+1$ new tokens (where $k$ is the maximum number of sentence tokens per example [^3] We follow Lee et al. [27] by using $k=20$ sentence tokens.):
@@
We augment our tokenizer function $t$ to form $t^{\prime}$, which maps sequence $s$ to tokens in $V^{\prime}$. In each sentence, we prepend a <SENT_i> token and append a <EOSEN> token as follows:
@@
The integer $i$ within <SENT_i> represents a sentence’s index, sampled uniformly at random from ${1,...,k}$. We sample sentence token IDs without replacement to eliminate information about the sentence’s position. We truncate sentences beyond the $k$’th to limit our vocabulary size.
Unlike the original SLM model (an encoder-only LM with an auxiliary pointer-decoder for pre-training), DEPTH is an encoder-decoder and cannot directly “point” to the latent vector of the next sentence. We address this by explicitly introducing <SENT_i> and <EOSEN> tokens into the vocabulary for the model to predict their IDs. Notably, we introduced <EOSEN> to signal that the next token is some <SENT_i> or the end of the example (<EOS>). In essence, <EOSEN> acts as a “signal” token so that the model can dynamically adjust its attention mask in the decoder (i.e., “should it attend to all previous tokens, or just to previous sentence tokens?”).
Formally, we define a pre-tokenization function $f$, which maps $s$ to $s^{\prime}$, where $s^{\prime}$ includes <SENT_i> and <EOSEN> tokens: $f:s\rightarrow s^{\prime}$. The tokenized input for DEPTH is produced with the function $T$, such that $T(s)=t(f(s))=X$. In practice, we use the SentencePiece [25] tokenizer as our function $t$, and adjust it to support DEPTH’s additional pre-training objective.
@@ 2.2 Corruption
We apply a corruption process to each batch of tokenized sequences. We sample masked spans using a geometric distribution (as in Joshi et al. [23] and Raffel et al. [45]), parameterized with an average span length of $\lambda$ and a masking probability of $p$. Spans overlapping with sentence tokens (<EOSEN> or <SENT_i>) are ignored. Masked token spans are replaced with a single sentinel token <special_token_z>, where ($z$) is a uniformly sampled integer from ${0,\ldots,99}$. The missing token sequences appear after the corresponding sentinel token in the labels. Note that in the original T5 implementation, sentinel tokens appear in the same incrementally decreasing order in each example (<special_token_99> followed by <special_token_98>, etc…). We randomly sample sentinel token ID’s for DEPTH to eliminate hints about sentence positions from the sentinel tokens. For example, with the T5 scheme for span-masking, the presence of <special_token_99> indicates that the sentence in which it appears comes first, making sentence un-shuffling too easy.
@@
Given an input sequence of up to $k$ sentences, we randomly shuffle the order of sentences in the model input (but do not shuffle the order of tokens within a sentence). We then task the model with reconstructing the original order of the sentence tokens in the target. We shuffle all examples in a batch with probability $p$ (we use $p=0.5$, as in SLM). We use this objective to encourage DEPTH to attend to other sentences for discourse-level cues, rather than relying on local words and phrases. By disrupting the original sentence order, DEPTH encourages learning of (1) the complete meaning of individual sentences, independent of their surrounding context, and (2) representations that encode how sentences relate to each other (e.g., discourse markers, co-reference, entailment). We show the full DEPTH corruption process in figure 1.
@@ 2.3 Attention Masks
This scheme encourages the model to rely on sentence token representations in the encoder to predict the correct next sentence token in the target. Our approach encourages sentence tokens in the input to encode discourse information (about the relative position of the sentence) that is useful for predicting the next sentence token via cross attention. As a second-order effect, this encourages non-sentence tokens in the encoder to provide their respective sentence token with useful discourse-level information.
@@ 2.4 Loss Formulation
Given that $\hat{S}$ is the set of all sentence tokens within our target sequence, let $Y=\{y_{1,1},y_{1,2},\ldots y_{m,len(m)}\}$ be the unified target sequence with tokens that can be either part of the span-masking task (non-sentence tokens) or the un-shuffling task (sentence tokens). The prediction $\hat{Y}=\{\hat{y}_{1,1},\hat{y}_{1,2},\ldots\hat{y}_{m,len(m)}\}$ represents the model’s output for each token in $Y$.
The loss for the DEPTH model, considering both reconstruction (a.k.a., span-masking) and sentence un-shuffling tasks, is simply the cross-entropy loss, which is calculated as:
@@
In this formulation, the summation runs over all sentences in the input (up to $k$), and each token within each sentence (up to $len(i)$, where $i$ is the index of some sentence) in the target sequence $Y$, thus allowing us to decompose DEPTH’s effectiveness in learning each pre-training task. We explore additional loss formulations in Appendix B.1.
@@ 3 Experimental Setup
The aim of our experiments is to measure the effectiveness of DEPTH against a standard encoder-decoder model. Accordingly, our experiments explore the learning dynamics of DEPTH model relative to a T5-Base (220M parameter) baseline. We pre-train both models on the C4 dataset [45, 14] to resemble the manner in which the original T5 was trained (see additional reasoning in Appendix D.2). We chose to use Base-sized models given limited computational resources, and ease of reproducibility (following the example of SLM, PMI, PEGASUS, and T5 [27, 28, 67, 45]). We did not use SLM as our baseline since its codebase and checkpoints are not released, and it cannot perform free-form text generation.
We chose to run our experiments without example packing since this is how the SLM model, which inspired DEPTH, was trained. Furthermore, example-packing in T5 enables unrelated examples to impact the model’s decisions, thereby harming performance [24, 49]. While example-packing is critical for more computationally-efficient training [13], we were interested in measuring which model was able to use training examples more effectively. We examine the impacts of avoiding example packing in Appendix A.2.
[…]
We conduct our pre-training experiments under two categories:
@@
    Continuous Pre-Training (CPT): Both T5 and DEPTH models are initialized from the T5-Base checkpoint on HuggingFace [63], and continue to pre-train on C4 with their respective objectives.
We note that our CPT experiments build on top of T5 models that have been trained for over 1T tokens, whereas the amount of tokens they see during continuous pre-training is relatively minuscule (67$\approx$ 67$\times$ 67 × fewer tokens for T5, and 80$\approx$ 80$\times$ 80 × fewer tokens for DEPTH). For reference, we show the relative pre-training configurations of similar models in Appendix D.1.

### v2

@@ Abstract
Language Models (LMs) struggle with linguistic understanding at the discourse level, even though discourse patterns such as coherence, cohesion, and narrative flow are prevalent in their pre-training data. To improve the discourse capabilities of LMs already at the pre-training stage, we introduce DEPTH, an encoder-decoder model that learns latent representations for sentences using a discourse-oriented pre-training objective. DEPTH combines hierarchical sentence representations with two objectives: (1) Sentence Un-Shuffling, and (2) Span-Corruption. Our approach trains the model to represent both sub-word-level and sentence-level dependencies over a pre-training corpora. When trained either from scratch or continuing from a pre-trained T5 checkpoint, DEPTH learns semantic and discourse-level representations faster than T5, outperforming it in span-corruption loss despite the additional sentence-un-shuffling objective. Evaluations on the GLUE, DiscoEval, and NI benchmarks demonstrate DEPTH’s ability to quickly learn diverse downstream tasks, which require syntactic, semantic, and discourse capabilities. Our approach extends the discourse capabilities of T5, while minimally impacting other natural language understanding (NLU) capabilities in the resulting LM. We share ur codebse for reproducibility: https://github.com/zbambergerNLP/depth.git.
DEPTH: Discourse Education through Pre-Training Hierarchically
Zachary Elisha Bamberger Technion zachary@campus.technion.ac.il                        Ofek Glick Technion ofek.glick@campus.technion.ac.il
Chaim Baskin Ben-Gurion University of the Negev chaimbaskin@bgu.ac.il                        Yonatan Belinkov Technion belinkov@technion.ac.il
@@ 1 Introduction
Discourse understanding—the ability to understand how sentences and broader textual units form cohesive narratives Miltsakaki et al. (2004); Prasad et al. (2008); Jernite et al. (2017); Prasad et al. (2018)—is fundamental to effective communication. However, LMs often struggle with this, especially when dealing with long and complex inputs, hindering their performance on tasks like persuasive argumentation (Durmus et al., 2019; Hidey et al., 2017; Chakrabarty et al., 2019), summarization (Zhao et al., 2022), essay scoring (Mim et al., 2021), dialogue systems (Hua et al., 2023), and following instructions (Wei et al., 2022a). Recent evidence from Yu et al. (2024) reinforces this view, demonstrating that human language comprehension occurs at multiple levels and that incorporating discourse-level objectives like next sentence prediction (NSP) can lead to more human-like language representations and improved contextual understanding.
Early attempts to incorporate discourse awareness into pre-training, such as Next Sentence Prediction (NSP) in BERT (Devlin et al., 2019) and Sentence Order Prediction (SOP) in ALBERT (Lan et al., 2020), proved overly simplistic and hindered learning effective discourse representations (Liu et al., 2019; Lan et al., 2020; Raffel et al., 2020). Subsequent encoder models like Sentence-level Language Model (SLM) (Lee et al., 2020), CONPONO (Iter et al., 2020), and Hi-Transformer (Wu et al., 2021) improved discourse capabilities in LMs but lacked generative capabilities.
Unlike the pre-training tasks for encoder LMs, next-token prediction provides decoder LMs like GPT (Radford et al., 2018, 2019; Brown et al., 2020; OpenAI et al., 2023) with powerful generative capabilities. However, without a dedicated and costly alignment phase (Ouyang et al., 2022; Wei et al., 2022a; Bai et al., 2022), these LMs tend to falter in understanding and executing human queries.
Even with a dedicated alignment phase, large decoder-only models generally perform poorly on discourse-oriented benchmarks that measure coherence and cohesion (Chen et al., 2019; Maimon and Tsarfaty, 2023a, b; Wang et al., 2023). Encoder-decoder models such as T5 (Raffel et al., 2020) and BART (Lewis et al., 2020a) consistently outperform much larger (400$\approx$ 400$\times$ 400 ×) decoder-only models like GPT-3 (Brown et al., 2020) and GPT-4 OpenAI et al. (2023) on these tasks. Recent work by Katz et al. (2024) indicates that incorporating encoder-decoder attention mechanisms into modern decoder-only models like Llama-3 Grattafiori et al. (2024), Qwen-2.5 Qwen et al. (2025), and Mistral Jiang et al. (2023) also boosts performance, suggesting potential additional benefits from pre-training models with such attention schemes.
@@
To improve the discourse-capabilities of encoder-decoders already at the pre-training stage, we propose DEPTH (Discourse-Education through Pre-Training Hierarchically), a hierarchical language model that learns representations for both sub-word and sentence-level tokens. DEPTH extends the pre-training objective of SLM from encoder-only models like BERT, to encoder-decoder models like T5. Notably, DEPTH introduces latent, heirarchical representations for sentences (as in Lee et al. (2020); Yang et al. (2021); Yu et al. (2024)) directly into the objective of a generative pre-training task. By employing a hierarchical attention mechanism across sub-word and sentence level tokens, DEPTH captures both fine-grained semantic dependencies and broader inter-sentence relationships. When pre-trained from scratch, our DEPTH model obtains meaningful representations for downstream tasks much faster than our baseline T5. Continuously pre-training T5 models with the DEPTH objective improves the discourse capabilities of the resulting models, without sacrificing performance in downstream NLU tasks.
@@ 2 Method
Pre-training DEPTH involves simultaneously performing span corruption (Raffel et al., 2020), while also un-shuffling sentences as in Lee et al. (2020). In Section 2.1 we introduce a new tokenizer for DEPTH, which combines T5’s tokenizer with the sentence segmentation operation required for sentence un-shuffling. In Section 2.2, we detail how we combined the pre-training objectives of both models into the sequence-to-sequence framework of T5. Next, in Section 2.3, we discuss how we introduce hierarchical representations during pre-training, and how this hierarchy encourages the model to learn discourse representations. Finally, in Section 2.4, we demonstrate how to combine the losses of T5 and SLM into a unified objective that is conducive to traditional teacher-forcing.
@@ 2.1 Tokenization
We introduce a tokenization function, $t$, which transforms an input string, $s$, into a sequence of tokens, $X=(x_{1,1},x_{1,2},\ldots,x_{m,\text{len}(m)})$, in our vocabulary, $V$. Each token $x_{i,j}$ denotes the $j$’th token of the $i$’th sentence, where $m$ is the number of sentences and $\text{len}(i)$ is the length of the $i$’th sentence. $V$ includes special tokens <EOS>, <BOS>, <PAD>, and sentinel tokens $V_{\text{sentinel}}$ of the form <special_token_z> as in the original T5 paper.
Figure 1: DEPTH tokenization and corruption process. Given an input document, DEPTH introduces sentence tokens (<SENT_i> and <EOSEN>), applies span masking, and shuffles sentences with probability 0.5. Attention patterns are shown with arrows (dotted for cross-attention, solid for self-attention).
To facilitate DEPTH’s hierarchical structure, we segment sentences using NLTK (Bird and Loper, 2004) and create $k+1$ new tokens [^1] We follow Lee et al. (2020), using $k=20$ sentence tokens.:
@@
$V^{\prime}$ $=V\cup S$
We augment our tokenizer function $t$ to form $t^{\prime}$, which maps sequence $s$ to tokens in $V^{\prime}$. In each sentence, we prepend a <SENT_i> token and append a <EOSEN> token:
@@
[…]
The integer $i$ in <SENT_i> represents a sentence’s index, sampled uniformly at random from $\{1,\ldots,k\}$ without replacement. We truncate sentences beyond the $k$’th to limit vocabulary size.
Unlike SLM (an encoder-only LM with an auxiliary pointer-decoder), DEPTH is an encoder-decoder that predicts <SENT_i> and <EOSEN> token IDs directly. The <EOSEN> token signals the next token is either <SENT_i> or <EOS>, allowing for dynamic attention masking in the decoder.
Formally, we define a pre-tokenization function $f:s\rightarrow s^{\prime}$, where $s^{\prime}$ includes <SENT_i> and <EOSEN> tokens. The tokenized input for DEPTH is produced with $T(s)=t(f(s))=X$. We use the SentencePiece (Kudo and Richardson, 2018) tokenizer as $t$, adjusted to support DEPTH’s sentence-level pre-training objective.
@@ 2.2 Corruption
We apply a corruption process to each batch of tokenized sequences. We sample masked spans using a geometric distribution (as in Joshi et al. (2020) and Raffel et al. (2020)), parameterized with an average span length of $\lambda$ and a masking probability of $p$. Spans overlapping with sentence tokens (<EOSEN> or <SENT_i>) are ignored. Masked token spans are replaced with a single sentinel token <special_token_z>, where ($z$) is a uniformly sampled integer from ${0,\ldots,99}$. The missing token sequences appear after the corresponding sentinel token in the labels. Note that in the original T5 implementation, sentinel tokens appear in the same incrementally decreasing order in each example (<special_token_99> followed by <special_token_98>, etc…). We randomly sample sentinel token ID’s for DEPTH to eliminate hints about sentence positions from the sentinel tokens. For example, with the T5 scheme for span-masking, the presence of <special_token_99> indicates that the sentence in which it appears comes first, making sentence un-shuffling too easy.
@@
Given an input sequence of up to $k$ sentences, we randomly shuffle the order of sentences in the model input [^2] We do not shuffle the order of tokens within a sentence. We then task the model with reconstructing the original order of the sentence tokens in the target. We shuffle all examples in a batch with probability $p=0.5$ (as in SLM). By disrupting the original sentence order, DEPTH encourages learning of (1) the complete meaning of individual sentences, independent of their surrounding context, and (2) representations that encode how sentences relate to each other [^3] E.g., discourse markers, co-reference, and entailment. We show DEPTH’s corruption process in Figure 1.
@@ 2.3 Attention masks
This scheme encourages the model to use sentence token representations in the encoder to predict the next sentence token in the decoder via cross-attention. It also ensures that non-sentence tokens in the encoder provide relevant discourse information to their corresponding sentence tokens.
@@ 2.4 Loss Formulation
Let $Y=\{y_{1,1},y_{1,2},\ldots,y_{m,\text{len}(m)}\}$ be the target sequence, where each token $y_{i,j}$ belongs either to the span-masking task (non-sentence tokens) or to the sentence un-shuffling task (sentence tokens). We denote by $\hat{S}$ the set of all sentence tokens in $Y$. The model prediction is given by $\hat{Y}=\{\hat{y}_{1,1},\hat{y}_{1,2},\ldots,\hat{y}_{m,\text{len}(m)}\}$, where $\hat{y}_{i,j}$ represents the predicted probability distribution over the vocabulary for token $y_{i,j}$. Let the total number of tokens be $N=|Y|$.
The loss for DEPTH, which jointly optimizes the reconstruction (span-masking) and sentence un-shuffling tasks, is defined as the token-averaged cross-entropy:
@@
In this formulation, the summation runs over all sentences in the input (i.e., up to $m$ sentences, where $1\leq m\leq k$), and within each sentence over its tokens. This allows us to decompose the model’s performance into the contributions from the sentence un-shuffling and the span-masking tasks. We explore additional loss formulations and weighting schemes in Appendix B.1.
@@ 3 Experimental setup
The aim of our experiments is to measure the effectiveness of DEPTH against a standard encoder-decoder model. Accordingly, our experiments explore the learning dynamics of DEPTH model relative to a T5-Base (220M parameter) baseline. We pre-train both models on the C4 dataset (Raffel et al., 2020; Dodge et al., 2021) to resemble the manner in which the original T5 was trained (see additional reasoning in Appendix D.2). We chose to use Base-sized models given limited computational resources, and ease of reproducibility (following the example of Lee et al. (2020); Levine et al. (2020); Zhang et al. (2020); Raffel et al. (2020)). We did not use SLM as our baseline since its codebase and checkpoints are not released, and it cannot perform free-form text generation.
We chose to run our experiments without example packing since this is how the SLM model, which inspired DEPTH, was trained. Furthermore, example-packing in T5 enables unrelated examples to impact the model’s decisions, thereby harming performance (Krell et al., 2021; Shi et al., 2024). While example-packing is critical for more computationally-efficient training (Ding et al., 2024), we were interested in measuring which model was able to use training examples more effectively. We examine the impacts of avoiding example packing in Appendix A.2.
Consistent with Nawrot (2023), we found that the Adafactor optimizer (Shazeer and Stern, 2018), while more computationally efficient, slightly harmed model performance. We therefore use AdamW (Loshchilov and Hutter, 2019) instead. We use a linearly increasing learning rate during the first $10,000$ steps, and then reduce the learning rate linearly for DEPTH (as in SLM), and with an inverse square learning rate ratio for T5 (as in the original T5 paper). We chose to use a masking probability of $p=0.3$, [^4] Raffel et al. (2020) reports that this span corruption ratio does not adversely impact downstream performance, although recently Ankner et al. (2024) suggested a dynamic masking rate tends to perform best. and an average span length of $\lambda=3$. Our mask probability is higher than the advised $0.15$ from T5 to accommodate for the fact that sentence-tokens within DEPTH cannot be masked.
We conduct two types of pre-training experiments:
@@
    Continuous Pre-Training (CPT): Both T5 and DEPTH models are initialized from the T5-Base checkpoint on HuggingFace (Wolf et al., 2019), and continue to pre-train on C4 with their respective objectives.
We note that our CPT experiments build on top of T5 models that have been trained for over 1T tokens, whereas the amount of tokens they see during continuous pre-training is relatively minuscule (67$\approx$ 67$\times$ 67 × fewer tokens for T5, and 80$\approx$ 80$\times$ 80 × fewer tokens for DEPTH). We compare configurations of similar-sized models in Appendix D.1.